OpenStudy (anonymous):

Find the intercepts to y=8-4x

OpenStudy (anonymous):

solve for the y-intercept by plugging in 0 to x > y=8-4(0) solve for y. then you can find the x-intercept by plugging in 0 to y (0)=8-4x solve for x and there you go

OpenStudy (anonymous):

if you have ever heard of y=mx+b m would be the slope and b would be the y intercept so 8 would be the intercept